Tom is three times as old as Mark. Twice Mark's age plus Tom's age is 90. How old is each person?
Answer by jorel1380(3719)

Tom=3Mark
2Mark+Tom=90.
-----------
Mark=1/3Tom
2/3Tom+Tom=90
5/3Tom=90
5Tom=270
Tom=54
Mark=18.